# Iterative Reasoning Preference Optimization

### v1

@@ Abstract
Iterative preference optimization methods have recently been shown to perform well for general instruction tuning tasks, but typically make little improvement on reasoning tasks (Yuan et al., 2024; Chen et al., 2024). In this work we develop an iterative approach that optimizes the preference between competing generated Chain-of-Thought (CoT) candidates by optimizing for winning vs. losing reasoning steps that lead to the correct answer. We train using a modified DPO loss (Rafailov et al., 2023) with an additional negative log-likelihood term, which we find to be crucial. We show reasoning improves across repeated iterations of this scheme. While only relying on examples in the training set, our approach results in increasing accuracy for Llama-2-70B-Chat from 55.6% to 81.6% on GSM8K (and 88.7% with majority voting out of 32 samples), from 12.5% to 20.8% on MATH, and from 77.8% to 86.7% on ARC-Challenge, which outperforms other Llama-2-based models not relying on additionally sourced datasets.
@@ 1 Introduction
Preference optimization has proven to give large gains when aligning pre-trained language models to human requirements compared to supervised fine-tuning alone (Ziegler et al., 2019; Stiennon et al., 2020). Offline methods such as DPO (Rafailov et al., 2023) are becoming more popular for their simplicity and efficiency. Recent results have shown that iterative application of such an offline procedure is beneficial, whereby the updated model is used to construct new preference relations that are more informative, and hence improve results further. These methods include Iterative DPO (Xu et al., 2023; Xiong et al., 2023), Self-Rewarding LLMs (Yuan et al., 2024), SPIN (Chen et al., 2024), and other methods (Rosset et al., 2024). Common to these approaches is that they have been shown to perform well on general instruction tuning tasks, but they either make only moderate gains or even decrease the performance on standard reasoning tasks. While other kinds of iterative training methods have been applied successfully to reasoning, particularly involving iteration of supervised fine-tuning (SFT) such as STaR (Zelikman et al., 2022) , RestEM (Singh et al., 2023), and V-STaR (Hosseini et al., 2024) [^1] V-STaR does use preference optimization, but for training a separate verifier model., using preference optimization to train the generative reasoning model is not applied in these methods.
@@ 2 Iterative Reasoning Preference Optimization
We assume we are given an initial model $M_{0}$, and a training set $D=\{x_{i},y_{i}\}$ containing questions $x_{i}$ and their correct answers $y_{i}$ The model will be trained and updated at each iteration, resulting in models $M_{0},M_{1},\dots M_{T}$.
@@
Next we build a dataset of preference pairs by selecting a winner response $(c_{i}^{w},y_{i}^{w})$ from $G_{i}^{w}$, and a loser response $(c_{i}^{l},y_{i}^{l})$ from $G_{i}^{l}$. In particular, we simply iterate over $G_{i}^{w}$ and $G_{i}^{l}$ simultaneously [^2] If the iteration reaches the end of a set, it restarts from the first element. to produce $K$ pairs $\{w_{k},l_{k}\}$, in order to ensure we use as much of the data as possible.
@@
$\mathcal{L}_{\text{DPO+NLL}}$ $=\mathcal{L}_{\text{NLL}}(x_{i},c_{i}^{w},y_{i}^{w})+\alpha\mathcal{L}_{\text{DPO}}(c_{i}^{w},y_{i}^{w},c_{i}^{l},y_{i}^{l}|x_{i})$
$=-\frac{\log M_{\theta}(x_{i},c_{i}^{w},y_{i}^{w})}{|x_{i}|+|c_{i}^{w}|+|y_{i}^{w}|}-\alpha\log\sigma\left(\beta\frac{\log M_{\theta}(c_{i}^{w},y_{i}^{w}|x_{i})}{\log M_{\text{t}}(c_{i}^{w},y_{i}^{w}|x_{i})}-\beta\frac{\log M_{\theta}(c_{i}^{l},y_{i}^{l}|x_{i})}{\log M_{\text{t}}(c_{i}^{l},y_{i}^{l}|x_{i})}\right).$ (1)
Here $M(x)$ denotes the probability of sequence $x$ under the model $M$, and $\sigma$ is the sigmoid function. We use the previous iteration’s model $M_{t}$ as the reference model in the denominator of the DPO term. Note that the NLL term is normalized by the total sequence length. The hyperparameter $\alpha$ balances the two loss terms. For brevity we omitted the pair index $k$, but we optimize this loss on each of the $k\in[1,K]$ pairs generated for every input sample. At the end of this training, we thus obtain our next model $M_{t+1}=M_{\theta}$, which will be then used to build data for the subsequent iteration.
@@
In our experiments, we define the models, and the training data they use as follows:
@@
This approach can be seen as a similar, but simpler, instance of the Self-Rewarding LLM training scheme proposed in Yuan et al. (2024), with three differences. Firstly, on each iteration in Self-Rewarding a new set of prompts is created to explore the input distribution, but in our approach we use the same fixed set of prompts. Secondly, due to this choice our experimental setup does not require a sophisticated reward model to judge the model generations, as we assume the training prompts have provided gold labels which we compare to. These two omitted steps are challenging for reasoning tasks because they require a language model to verify correctness, which is known to be difficult (Huang et al., 2023). Thirdly, we show that our DPO+NLL objective is important for our reasoning tasks, whereas Self-Rewarding LLM’s used the standard DPO objective.
@@ 3.1 Math word problems: GSM8K
As a seed model $M_{0}$ we use the chat version of Llama-2 70B model (Touvron et al., 2023), which is instruction finetuned. We use a zero-shot prompt containing the question together with instructions to produce a chain-of-thought and to follow a specific format so the final answer can be easily extracted (the exact prompt is given in subsection A.1). In each iteration, we generate $N=30$ solutions per problem using sampling with temperature 0.8 for iterations 1–2 and temperature 1.3 for iterations 3–4 (hoping that there is a significant number of incorrect generations in later iterations). Since some problems might not have any model-generated correct solution, we include the gold human written solution $(c_{i},y_{i})$ in the winning set $G_{i}^{w}$ so it is not empty. Then we generate $K=10$ pairs per problem for training with our loss in Equation 1, and filtered out examples that were too long in terms of overflowing the context length or else do not have any incorrect generations. This gave around 55–60k pairs for training, per iteration.
In total, we performed 4 iterations, producing models $M_{1}$, $M_{2}$, $M_{3}$ and $M_{4}$. For each iteration, we train a maximum of 5000 steps, then select the best checkpoint using a held-out 1k samples from the training set. We then retrain including those 1k samples for the selected number of steps. The coefficient $\alpha$ is tuned in {0.5, 1, 2, 4} when training M1, and we end up using 1 for all experiments in the paper. We used a batch size of 16 and a learning rate 7e-7.
@@
We find that Iterative RPO outperforms zero-shot CoT, supervised finetuning (SFT) on the gold CoT solutions and variants of DPO by a wide margin. SFT gives a boost in performance compared to zero-shot CoT from 55.6% to 63.5%, but this is still far from the 81.6% of Iterative RPO. We apply standard DPO to the same set of preference pairs $D_{0}^{\text{pairs}}$ as used in the first iteration of our method. Whether initializing from Llama-2-70b-chat ($M_{0}$) or from SFT training on the chosen (winner) examples, we find that DPO performance, while being better than zero-shot CoT, is no better than the SFT model, with accuracies of 61.8% or 60.3% respectively. We also show that SFT on only the chosen CoT solutions, which corresponds to the first iteration of the STaR method, improves results to 65.2% over SFT on the gold solutions alone, but still falls short of the performance of the first iteration of Iterative RPO. One hypothesis for these improvements is the necessity of including the rejected sequences in the training objective, otherwise their probability increases along with the chosen samples, see Figure 3. We note this observation has also been reported in concurrent work (Hong et al., 2024). All of the results reported above are using a single generation at test time using greedy decoding. If we use majority voting over 32 samples (sampling with temperature 0.8), a standard approach to improve performance in the literature, we can improve the accuracy of our approach from 81.1% to 88.2% for iteration 3, and from 81.6% to 88.7% for iteration 4 of Iterative RPO.
Iterations of Iterative RPO yield improve reasoning
@@
We observe that Iterative RPO provides improvements over its training iterations, increasing the base model accuracy by 47% (from 55.6% to 81.6%) in total. In contrast, supervised training using the gold CoT only brings about a 14% accuracy boost. We see performance improves across each iteration, from 73.1% to 78.0% to 81.1% to 81.6%. However, the gain decays across the iterations (17.5%, 4.9%, 3.1%, 0.5%), indicating an upper limit on learning across iterations, especially as we are iterating across a fixed number of prompts, i.e., only from the training samples. We also show that it is the iterations of updating the model (i.e., initializing from the previous model) that are helping, not just because there is more data in the form of new pairs generated from the fixed training set. To test this we run the first iteration of Iterative RPO but on twice as much paired data, as well as the STaR method first iteration with twice as much data as well. In both cases performance improves compared to less data, but not as much as performing two iterations. Iterative RPO with twice as much data obtains 74.8% (an improvement over 73.1% using the original dataset size); however, training for two iterations obtains 78.0%. For STaR, training on twice as much data obtains 66.9%, compared to 65.2% with the original data, which is still a much lower performance than Iterative RPO.
@@
The first iteration of our method can be compared to standard DPO training, which uses the same preference data, as reported in Table 1. We see a large performance drop (73.1% vs. 61.8%) using DPO compared to our method after 1 iteration. The gap remains large even when the standard DPO training starts from the superior SFT-tuned model, which it has been argued improves DPO’s performance (Rafailov et al., 2023, 2024). Our results support the need of the NLL loss term in our training, not just using SFT for initialization. To further understand this, we plot the sequence-level log probability over training steps for these methods in Figure 2. We see that for DPO without NLL loss there is a decrease over training for the chosen sequences, whereas for DPO with NLL there is not, which may help explain the improved performance of the latter. We note that related observations have been made elsewhere in various settings (Pal et al., 2024; Xu et al., 2024). Further, we note that whether we initialize with Llama-2-70b-chat or SFT on chosen for Iterative RPO, accuracy results of first iteration training do not seem to deviate (both obtain the same score 73.1%).
Figure 3: Effect of SFT training: although SFT training (solid green) is on chosen sequences only, the rejected sequence probability also increases and is close to the chosen sequence probability. We show the log probability vs. number of training steps for SFT training (on only the chosen examples from $D_{0}^{\text{pairs}}$) on GSM8K. The solid curve shows the log probabilities of sequences used for training. The dotted curve shows the log probabilities of the rejected sequences. Although those sequences are not used for SFT training, the log probabilities of those lower-quality sequences also increase. In contrast, our DPO+NLL training (blue) manages to decrease the rejected probabilities while increasing the chosen probabilities. This observation could potentially help explain why SFT-only performance lags significantly behind Iterative RPO Iteration 1 performance.
@@ 3.2 ARC-Challenge task
To test reasoning capabilities outside of mathematics, we employ ARC (Clark et al., 2018) which covers multiple science subjects. The dataset contains 7.7k multiple-choice questions split into easy and challenge sets. We report results on the ARC-Challenge test set which has 1172 examples. There is no gold chain-of-thought reasoning provided for training examples in this task, which in any case is not required in our method, as we only compute rewards based on the final answer. One consequence is that if there is no model-generated correct solution for a question, then that question is not included in our training. We thus follow the same setup as before to first generate reasoning and then a final answer by the models (see subsection A.1 for prompt) to construct data for iterations of Iterative RPO. We only train on the training set and do not utilize ARC Corpus.
@@ 3.3 MATH task
In each iteration, we generate $N=20$ solutions per problem using sampling with temperature 0.8 for iterations 1–2 and temperature 1.0 for iteration 3. We select $K=15$ pairs of solutions per problem, and after filtering out pairs with overly long generations, for each iteration we randomly select around 75k example pairs. We train a maximum of 5000 steps per iteration; other details are similar to GSM8K setups.
Results are given in Table 2. We again find that Iterative RPO provides increased performance across iterations, from 17.7% to 19.9% to 20.8% over three iterations. These results outperform few-shot CoT (12.5%), SFT on chosen sequences (16.8%) and standard DPO (12.4%). In particular, DPO degrades the performance compared to initialization.
@@ 4 Related Work
Several works have implemented iterative reinforcement learning from human feedback (RLHF) with a human-in-the-loop to provide additional labels to retrain the reward model at each iteration, e.g., via Proximal Policy Optimization (PPO) (Schulman et al., 2017), reporting improvements across iterations (Bai et al., 2022; Touvron et al., 2023). Recently, approaches have been proposed to perform iterative alignment without a human-in-the-loop. Iterative DPO (Xu et al., 2023; Xiong et al., 2023) optimizes preference pairs using DPO (Rafailov et al., 2023) at each iteration, and then constructs new preference pairs for the next iteration by generating them using the updated model, and scoring them using a reward model. Other iterative methods than DPO exist as well, such as the Cringe loss (Adolphs et al., 2023), Pairwise Cringe Loss (Xu et al., 2023) and ReST (Gulcehre et al., 2023).
@@
Self-Rewarding LLMs (Yuan et al., 2024) also use Iterative DPO with the LLM itself used as a reward model to construct pairs for each successive iteration. Both that work, and the work of Rosset et al. (2024) and Snorkel AI Team (2023) which do similar iterations but with external reward models, show significant gains on general instruction following tasks. However, again, only modest gains on reasoning tasks are reported.
@@
While a number of approaches have been developed to curate or distill training data for reasoning tasks (Yu et al., 2023; Toshniwal et al., 2024), in this work we focus on learning algorithms which is an orthogonal axis. Expert Iteration assumes a reward model, and repeatedly uses rejection sampling to filter generations and train on them, which is found to match the sample complexity of PPO (Havrilla et al., 2024). STaR (Zelikman et al., 2022) relies on a similar loop: generate rationales to answer many questions, prompted with a few rationale examples; if the generated answers are wrong, try again to generate a rationale given the correct answer; and then fine-tune on all the rationales that ultimately yielded correct answers; and repeat. ReSTEM (Singh et al., 2023) assumes a ground truth verifier and also fine-tunes on filtered samples in a repeated fashion. All these methods rely on finding high-quality samples for SFT-like training, rather than using DPO-like pairwise preference optimization as in our work.

### v2

@@ Abstract
Iterative preference optimization methods have recently been shown to perform well for general instruction tuning tasks, but typically make little improvement on reasoning tasks (Yuan et al., 2024; Chen et al., 2024). In this work we develop an iterative approach that optimizes the preference between competing generated Chain-of-Thought (CoT) candidates by optimizing for winning vs. losing reasoning steps that lead to the correct answer. We train using a modified DPO loss (Rafailov et al., 2023) with an additional negative log-likelihood term, which we find to be crucial. We show reasoning improves across repeated iterations of this scheme. While only relying on examples in the training set, our approach results in increasing accuracy on GSM8K, MATH, and ARC-Challenge for Llama-2-70B-Chat, outperforming other Llama-2-based models not relying on additionally sourced datasets. For example, we see a large improvement from 55.6% to 81.6% on GSM8K and an accuracy of 88.7% with majority voting out of 32 samples.
@@ 1 Introduction
Preference optimization has proven to give large gains when aligning pre-trained language models to human requirements compared to supervised fine-tuning alone (Ziegler et al., 2019; Stiennon et al., 2020). Offline methods such as DPO (Rafailov et al., 2023) are becoming more popular for their simplicity and efficiency. Recent results have shown that iterative application of such an offline procedure is beneficial, whereby the updated model is used to construct new preference relations that are more informative, and hence improve results further. These methods include Iterative DPO (Xu et al., 2023; Xiong et al., 2023), Self-Rewarding LLMs (Yuan et al., 2024), SPIN (Chen et al., 2024), and other methods (Rosset et al., 2024). Common to these approaches is that they have been shown to perform well on general instruction tuning tasks, but they either make only moderate gains or even decrease the performance on standard reasoning tasks. While other kinds of iterative training methods have been applied successfully to reasoning, particularly involving iteration of supervised fine-tuning (SFT) such as STaR (Zelikman et al., 2022), RestEM (Singh et al., 2024), and V-STaR (Hosseini et al., 2024) [^1] V-STaR does use preference optimization, but for training a separate verifier model., using preference optimization to train the generative reasoning model is not applied in these methods.
@@ 2 Iterative Reasoning Preference Optimization
We assume we are given an initial model $M_{0}$, and a training set $D=\{x_{i},y_{i}\}$ containing questions $x_{i}$ and their correct answers $y_{i}$ The model will be trained and updated at each iteration, resulting in models $M_{0},M_{1},\dots,M_{T}$.
@@
Next we build a dataset of preference pairs by selecting a winner response $(c_{i}^{w},y_{i}^{w})$ from $G_{i}^{w}$, and a loser response $(c_{i}^{l},y_{i}^{l})$ from $G_{i}^{l}$. In particular, we simply iterate over $G_{i}^{w}$ and $G_{i}^{l}$ simultaneously [^2] If the iteration reaches the end of a set, it restarts from the first element. If one of the sets is empty, then that input will be ignored. to produce $K$ pairs $\{w_{k},l_{k}\}$, in order to ensure we use as much of the data as possible.
@@
$\mathcal{L}_{\text{DPO+NLL}}$ $=\mathcal{L}_{\text{NLL}}(c_{i}^{w},y_{i}^{w}|x_{i})+\alpha\mathcal{L}_{\text{DPO}}(c_{i}^{w},y_{i}^{w},c_{i}^{l},y_{i}^{l}|x_{i})$
$=-\frac{\log M_{\theta}(c_{i}^{w},y_{i}^{w}|x_{i})}{|c_{i}^{w}|+|y_{i}^{w}|}-\alpha\log\sigma\left(\beta\frac{\log M_{\theta}(c_{i}^{w},y_{i}^{w}|x_{i})}{\log M_{t}(c_{i}^{w},y_{i}^{w}|x_{i})}-\beta\frac{\log M_{\theta}(c_{i}^{l},y_{i}^{l}|x_{i})}{\log M_{t}(c_{i}^{l},y_{i}^{l}|x_{i})}\right).$ (1)
Here $M(x)$ denotes the probability of sequence $x$ under the model $M$, and $\sigma$ is the sigmoid function. We use the previous iteration’s model $M_{t}$ as the reference model in the denominator of the DPO term. Note that the NLL term is normalized by the total response length. The hyperparameter $\alpha$ balances the two loss terms. For brevity we omit the pair index $k$, but we optimize this loss on each of the $k\in[1,K]$ pairs generated for every input sample. At the end of this training, we thus obtain our next model $M_{t+1}=M_{\theta}$, which will be then used to build data for the subsequent iteration.
@@
In our experiments, we define the models and the training data they use as follows:
@@
This approach can be seen as a similar, but simpler, instance of the Self-Rewarding LLM training scheme proposed in Yuan et al. (2024), with three differences. Firstly, on each iteration in Self-Rewarding a new set of prompts is created to explore the input distribution, but in our approach we use the same fixed set of prompts. Secondly, due to this choice our experimental setup does not require a sophisticated reward model to judge the model generations, as we assume the training prompts have provided gold labels which we compare to. These two omitted steps are challenging for reasoning tasks because they require a language model to verify correctness, which is known to be difficult (Huang et al., 2024). Thirdly, we show that our DPO+NLL objective is important for our reasoning tasks, whereas Self-Rewarding LLM has used the standard DPO objective.
@@ 3.1 Math Word Problems: GSM8K
As a seed model $M_{0}$ we use the chat version of Llama-2 70B model (Touvron et al., 2023), which is instruction fine-tuned. We use a zero-shot prompt containing the question together with instructions to produce a chain-of-thought and to follow a specific format so the final answer can be easily extracted (the exact prompt is given in subsection A.1). In each iteration, we generate $N=30$ solutions per problem using sampling with temperature 0.8 for iterations 1–2 and temperature 1.3 for iterations 3–4 (hoping that there is a significant number of incorrect generations in later iterations). Since some problems might not have any model-generated correct solution, we include the gold human written solution $(c_{i},y_{i})$ in the winning set $G_{i}^{w}$ so it is not empty. Then we generate $K=10$ pairs per problem for training with our loss in Equation 1, and filtered out examples that were too long in terms of overflowing the context length or else do not have any incorrect generations. This gave around 55–60k pairs for training, per iteration. [^3] If after filtering the number of pairs is larger than 60k, then we randomly select around 60k examples. This number is fixed because we do not want to introduce another source of variability in our experiments.
In total, we performed four iterations, producing models $M_{1}$, $M_{2}$, $M_{3}$, and $M_{4}$. For each iteration, we train a maximum of 5000 steps, then select the best checkpoint using a held-out 1k samples from the training set. We then retrain including those 1k samples for the selected number of steps. The coefficient $\alpha$ is tuned in {0.5, 1, 2, 4} when training M1, and we end up using 1 for all experiments in the paper. The coefficient $\beta$ in the DPO loss is tuned in {0.05, 0.1, 0.5, 1.0}, and we end up using 0.1 in this experiment. We used a batch size of 16 and a learning rate 7e-7.
@@
We find that Iterative RPO outperforms zero-shot CoT, supervised fine-tuning (SFT) on the gold CoT solutions, and variants of DPO by a wide margin. SFT gives a boost in performance compared to zero-shot CoT from 55.6% to 63.5% but still far from the 81.6% of Iterative RPO. We apply standard DPO to the same set of preference pairs $D_{0}^{\text{pairs}}$ as used in the first iteration of our method. Whether initializing from Llama-2-70b-chat ($M_{0}$) or from SFT training on the chosen (winner) examples, we find that DPO performance, while being better than zero-shot CoT, is no better than the SFT model, with accuracies of 61.8% or 60.3% respectively. We also show that SFT on only the chosen CoT solutions, which corresponds to the first iteration of the STaR method, improves results to 65.2% over SFT on the gold solutions alone, but still falls short of the performance of the first iteration of Iterative RPO. One hypothesis for these improvements is the necessity of including the rejected sequences in the training objective, otherwise their probability increases along with the chosen samples, see Figure 2. We note this observation has also been reported in concurrent work (Hong et al., 2024). All of the results reported above are using a single generation at test time using greedy decoding. If we use majority voting over 32 samples (sampling with temperature 0.8), a standard approach to improve performance in the literature, we can improve the accuracy of our approach from 81.1% to 88.2% for iteration 3, and from 81.6% to 88.7% for iteration 4 of Iterative RPO.
Iterations of Iterative RPO yield improved reasoning
@@
We observe that Iterative RPO provides improvements over its training iterations, increasing the base model accuracy by 47% (from 55.6% to 81.6%) in total. In contrast, supervised training using the gold CoT only brings about a 14% accuracy boost. We see performance improves across each iteration, from 73.1% to 78.0% to 81.1% to 81.6%. However, the gain decays across the iterations (17.5%, 4.9%, 3.1%, 0.5%), indicating an upper limit on learning across iterations, especially as we are iterating across a fixed number of prompts, i.e., only from the training samples. We also show that it is the iterations of updating the model (i.e., initializing from the previous model) that are helping, not just because there is more data in the form of new pairs generated from the fixed training set. To test this statement, we run the first iteration of Iterative RPO but on twice as much paired data by doubling $K$, as well as the STaR method first iteration with twice as much data as well. In both cases performance improves compared to less data, but not as much as performing two iterations. Iterative RPO with twice as much data obtains 74.8% (an improvement over 73.1% using the original dataset size); however, training for two iterations obtains 78.0%. For STaR, training on twice as much data obtains 66.9%, compared to 65.2% with the original data, which is still a much lower performance than Iterative RPO.
(a) SFT trained on chosen seqs; init from Llama
(b) SFT trained on gold CoTs; init from Llama
Figure 2: Effect of SFT training. (a) Although SFT training (solid green) is on chosen sequences ($D_{0}^{\text{pairs}}$, from iterative RPO iteration 1) only, the rejected sequence log probabilities also increase and are close to the chosen sequence probabilities. In contrast, our DPO+NLL training (blue) manages to decrease the rejected probabilities while increasing the chosen probabilities. This observation could potentially help explain why SFT-only performance lags significantly behind Iterative RPO Iteration 1 performance. (b) We show a similar plot but where SFT is trained on gold (dataset-provided) CoTs. Chosen and rejected sequence probabilities (which are from $D_{0}^{\text{pairs}}$) are still close to each other, but with a slightly bigger gap. In addition, the chosen sequence probabilities barely increase.
@@
The first iteration of our method can be compared to standard DPO training, which uses the same preference data, as reported in Table 1. We see a large performance drop (73.1% vs. 61.8%) using DPO compared to our method after 1 iteration. The gap remains large even when the standard DPO training starts from the superior SFT-tuned model, which it has been argued improves DPO’s performance (Rafailov et al., 2023, 2024). Our results support the need of the NLL loss term in our training, not just using SFT for initialization. To further understand this, we plot the sequence-level log probability over training steps for these methods in Figure 3. We see that for DPO without NLL loss there is a decrease over training for the chosen sequences, whereas for DPO with NLL there is not, which may help explain the improved performance of the latter. We note that related observations have been made elsewhere in various settings (Pal et al., 2024; Xu et al., 2024; Hong et al., 2024). Further, we note that whether we initialize with Llama-2-70b-chat or SFT on chosen for Iterative RPO, accuracy results of first iteration training do not seem to deviate (both obtain the same score 73.1%).
@@ 3.2 ARC-Challenge Task
To test reasoning capabilities outside of mathematics, we employ ARC (Clark et al., 2018) which covers multiple science subjects. The training dataset contains 7.7k multiple-choice questions split into easy and challenge sets. We report results on the ARC-Challenge test set which has 1172 examples. There is no gold chain-of-thought reasoning provided for training examples in this task, which in any case is not required in our method, as we only compute rewards based on the final answer. One consequence is that if there is no model-generated correct solution for a question, then that question is not included in our training. We follow the same setup as before to first generate reasoning and then a final answer by the models (see subsection A.1 for prompt) to construct data for iterations of Iterative RPO. We only train on the training set (both easy and challenge sets) and do not utilize the ARC Corpus.
@@ 3.3 MATH Task
In each iteration, we generate $N=20$ solutions per problem using sampling with temperature 0.8 for iterations 1–2 and temperature 1.0 for iteration 3. We select $K=15$ pairs of solutions per problem, and after filtering out pairs with overly long generations, for each iteration we end up with around 75k example pairs. We train a maximum of 5000 steps per iteration; other details are similar to GSM8K setups.
Results are given in Table 2. We again find that Iterative RPO provides increased performance across iterations, from 17.7% to 19.9% to 20.8% over three iterations. Majority voting (32 samples, temperature 0.8) leads to a significant boost in performance (29.1%). These results outperform few-shot CoT (12.5%), SFT on chosen sequences (16.8%) and standard DPO (12.4%). In particular, DPO degrades the performance compared to initialization.
@@ 4 Related Work
Several works have implemented iterative reinforcement learning from human feedback (RLHF) with a human-in-the-loop to provide additional labels to retrain the reward model at each iteration, e.g., via Proximal Policy Optimization (PPO) (Schulman et al., 2017), reporting improvements across iterations (Bai et al., 2022; Touvron et al., 2023). Recently, approaches have been proposed to perform iterative alignment without a human-in-the-loop. Iterative DPO (Xu et al., 2023; Xiong et al., 2023) optimizes preference pairs using DPO (Rafailov et al., 2023) at each iteration, and then constructs new preference pairs for the next iteration by generating them using the updated model, and scoring them using a reward model. Other iterative methods than DPO exist as well, such as the Cringe loss (Adolphs et al., 2023), Pairwise Cringe Loss (Xu et al., 2023), and ReST (Gulcehre et al., 2023).
@@
Self-Rewarding LLMs (Yuan et al., 2024) also use Iterative DPO with the LLM itself used as a reward model to construct pairs for each successive iteration. Both that work and the work of Rosset et al. (2024) and Snorkel AI Team (2023), which do similar iterations but with external reward models, show significant gains on general instruction following tasks. However, again, only modest gains on reasoning tasks are reported.
@@
While a number of approaches have been developed to curate or distill training data for reasoning tasks (Yu et al., 2023; Toshniwal et al., 2024), in this work we focus on learning algorithms which is an orthogonal axis. Expert Iteration assumes a reward model, and repeatedly uses rejection sampling to filter generations and train on them, which is found to match the sample complexity of PPO (Havrilla et al., 2024). STaR (Zelikman et al., 2022) relies on a similar loop: generate rationales to answer many questions, prompted with a few rationale examples; if the generated answers are wrong, try again to generate a rationale given the correct answer; and then fine-tune on all the rationales that ultimately yielded correct answers; and repeat. ReSTEM (Singh et al., 2024) assumes a ground truth verifier and also fine-tunes on filtered samples in a repeated fashion. All these methods rely on finding high-quality samples for SFT-like training, rather than using DPO-like pairwise preference optimization as in our work.